OpenStudy (anonymous):

which of the following is NOT a solution of 2a^2+3b=5? a) a=0, b=5/3 b) a=1, b=1 c) a=2, b=-1 d) a=3, b=-4 e) a=4, b=-9

OpenStudy (anonymous):

try plugging in the possible choices into the equation.. like this: a) a=0, b=5/3, so.. 2(0)^2+3(5/3)=5 2+5 = 5 [False]

OpenStudy (campbell_st):

lol... wow 2(0)^2 = 2.... perhaps it should be 2(0)^2 + 3(5/3) = 0 + 5 = 5 true

OpenStudy (anonymous):

but the answer is d though D: